Can you piease Show hou got the answer neps me Undlerstand H better hank you So much
1 6. A business is trying to identify the optimal order quantity for one of its produets. Demand 2 is constant and sells about 300 per month. Assume that the unit cost of the product is $30.00 and it costs $85 dollars to place an order. Annual holdings costs are 15% of the value 4 of the inventory. The lead times is 10 days. Answer the folloiwng inventory policy questions: 6 a. If the cost of the product is $20 per unit, what is the Eoo? 7 b. Assume 250 days of the operation per year. What is the reorder point? 8 c. If the lead time is increased to 20 days, what is the reorder point? 9 d. Comparing the value from part e with the EOQ for part a, explain the relative size of these two quantities 10 d. W hat is the reorder point for part (c) if the reorder point is expressed in terms of the 12 inventory on hand rather than the inventory position? 13 14 15
Can you piease Show hou got the answer neps me Undlerstand H better hank you So much
Answer #1

(a)

Annual demand, D = 300 x 12 = 3,600
Ordering cost, K = $85
Cost of the product, C = $20
Annual holding cost %, i = 15%

EOQ = (2.D.K / i.C)1/2 = sqrt(2*3600*85 / (0.15*20)) = 452 units

(b)

Average daily demand, d = D / 250 = 3600 / 250 = 14.4
Lead time, L = 10 days

So, Reorder point, ROP = lead time demand = d.L = 144 units

(c)

ROP = lead time demand = d.L = 14.4 x 20 = 288 units

(d)

The EOQ and ROP values completes the policy variables i.e. in each order cycle, 452 units are to be ordered when the inventory position falls to 288 units.

(e)

Inventory position = Inventory on-hand + quantity on order - backorder

In this case:

  • EOQ = 452, so, the time between orders = 452/300 = 1.5 months, and the lead time is only 10 days. So, there is no question of quantity on order while placing a new order.
  • Backorder is not allowed.

So, there will be no difference in inventory level and inventory position and hence the ROP will also be the same.